this is not a word problem just a fraction problem. the math problem is
2 5/8 TIMES 6 2/3 ( two and fife eights times 6 and two thirds )
i dont need an explanation, because i will be able to figure it out on my own. all i need is the answer because my math teacher is making me do it a certain way so it would be a little pointless for you to explain it! thank you very much! :)

Answer by stanbon(75887)

2 5/8 TIMES 6 2/3
--------
= (21/8)(20/3)
----
Reduce the fractions:
= (7/2)(5/1)
-----
= 35/2
----
= 17 1/2
=====================
Cheers,
Stan H.
======================